The sum of the digits of a two-digit number is 11. The value of the number is eight more than 10 times the tens digit. Find the number
pls... answer pls...
ADVANCE THANK YOU!!!

Answer by sudhanshu_kmr(1152)


let unit digit = x then tens digit = 11-x
value of number = 10(11-x) + x
10(11-x) + x = 8 + 10(11-x) according to question.....
=> 110 - 10x + x = 8 + 110 - 10x
=> x = 8
so, tens digit = 3
number is : 38